# Reproducible Machine Learning-based Voice Pathology Detection: Introducing the Pitch Difference Feature

### v2

@@ Abstract
This study introduces a novel methodology for voice pathology detection using the publicly available Saarbrücken Voice Database (SVD) database and a robust feature set combining commonly used acoustic handcrafted features with two novel ones: pitch difference (relative variation in fundamental frequency) and a NaN feature (failed fundamental frequency estimation).
We evaluate six machine learning (ML) classifiers — support vector machine, k-nearest neighbors, naive Bayes, decision tree, random forest, and AdaBoost — using grid search for feasible hyperparameters of selected classifiers and 20480 different feature subsets. Top 1000 classifier-feature subset combinations for each classifier type are validated with repeated stratified cross-validation. To address class imbalance, we apply K-Means SMOTE to augment the training data.
Our approach achieves outstanding performance, reaching 85.61%, 84.69% and 85.22% unweighted average recall (UAR) for females, males and combined results respectivelly. We intentionally omit accuracy as it is a highly biased metric for imbalanced data. This advancement demonstrates significant potential for clinical deployment of ML methods, offering a valuable supportive tool for an objective examination of voice pathologies. To enable an easier use of our methodology and to support our claims, we provide a publicly available GitHub repository with DOI 10.5281/zenodo.13771573. Finally, we provide a REFORMS checklist to enhance readability, reproducibility and justification of our approach.
Graphical abstract
{strip}
@@
Introducing reproducible machine-learning-based system for voice pathology detection. Novel pitch difference and NaN features improving classification performance. Addressing class imbalance with k-means SMOTE, boosting minority class predictions. Avoiding data leakage by managing multiple recordings per patient effectively. Performing extensive grid search for optimal features and hyperparameters.
@@ 1 Introduction
Voice disorders are characterized by the alteration of voice quality due to functional or morphological changes. Although these disorders are widespread, they are often underestimated, which can delay proper healing. Voice disorder diagnosis consists of several medical examinations, including acoustic analysis. This involves the processing and analysis of various acoustic features of the voice signal, which can reveal changes or alterations in voice quality caused by specific diseases [1]. In this context, artificial intelligence (AI) can be a valid and powerful tool as it may support the decision making process in clinical settings. However, utilization of these solutions imposes many additional tasks, ranging from data collection and preprocessing to ground truth labeling and correct algorithm selection.
@@
In this paper on voice pathology detection, we introduce a methodology for Saarbrücken Voice Database (SVD) that addresses the presence of multiple recordings from the same patients. Without appropriate handling, this can lead to data leakage and overly optimistic results. We consider this an important contribution, as this problem is rarely reflected in the literature. We also present two features, pitch difference and NaN feature, that reflect the fundamental frequency variability and experimentally verify their importance. Based on the experimental results, we highlight the importance of training classifiers separately for male and female patients. As we do not investigate the relationship between feature causality or correlation with pathology and we work with a large number of features, we employ machine learning (ML) algorithms as suitable classifiers. Our feature selection process revealed that the optimal feature sets may differ between sexes and classifiers. To identify robust feature sets, we train multiple classifiers with various hyperparameters for different combinations of feature subsets (see Table 2). By estimating the mean Matthews correlation coefficient (MCC), we identify the most promising combinations of feature sets and classifiers. Subsequently, we conduct repeated stratified 10-fold cross-validation on top performing combinations of classifiers and feature subsets to estimate the average performance metrics and their corresponding standard deviations in order to estimate the confidence in our results. Contrary to many studies, our approach encompasses all possible pathologies rather than focusing on a selected few, ensuring that our models reflect the diversity of pathologies present in the population.
The document is organized as follows. In Section 2, we explore the current state-of-the-art practices in the field of pathology detection from voice recordings. In Section 3, we describe the dataset that was used for the development of our feature set. In Section 4, we describe the first preprocessing step to obtain the feature set. In Section 5, we explain the augmentation technique used to address the class imbalance in the used dataset. In Section 6, we describe the algorithms used as classifiers for healthy and pathological patients. In Section 7, we introduce the metrics used to determine the effectiveness of the developed model, with a proper explanation of the chosen metrics. Lastly, in the Section 8, we interpret the results,compare them with the state-of-the-art models and summarize our findings in the Section 9.
Due to the complexity of our methodology, we provide a flowchart to show the process of feature extraction, classifier training and results validation in Figure 1.
Figure 1: Flow diagram of methodology
@@ 2 Related Works
In this chapter, we outline the original contribution related to voice pathology detection present in the literature, limiting our research to works connected to SVD. This choice is due to its comprehensiveness: it is the only voice recording database representing the common voice pathologies. See Section 3 for detailed information.
@@
In [4], authors of the study test various acoustic features, such as pitch, jitter, shimmer, harmonic-to-noise ratio (HNR), detrended fluctuation analysis parameters, glottis quotients (open, closed), glottal-to-noise excitation ratio, Teager–Kaiser energy operator, modulation energy, and normalized noise energy as well as mel-frequency cepstral coefficients (MFCC). Moreover, they considered the sound samples and their spectrograms as input, all in combination with XGBoost, IsolationForest, and DenseNet models to determine the pathologic samples, reaching an F1 score of 73.3% and unweighted average recall (UAR) we computed as 73.3%. While, in [5], features derived from self-supervised learning models, Data2Vec and Wav2Vec, alongside MFCC are explored. The reliability of these features to evaluate voice quality is tested using support vector machine (SVM) and deep neural networks (DNN), achieving an accuracy of 77.83% and UAR we computed as 77.86%.
@@
A different approach is taken in [6], where the authors transform sound wave data into spectrograms and treat classification as an image recognition problem using a convolutional neural network (CNN), achieving 73.93% accuracy and UAR we computed as 70.68%.
@@
In [7], the authors use MFCC and gammatone frequency cepstral coefficients (GFCC) with a classifier based on a neural network (NN), achieving 81.84% accuracy. Unlike other studies that use the /a:/ sound for feature extraction, they employ whole sentences. Unfortunately, we were not able to compute UAR, because the reported results appear inconsistent, making it impossible to calculate the metric. Another study [8] tests a DNN with convolutional layers, using the voice signal as a feature set for the convolutional layer, achieving 68.08% accuracy and UAR we computed as 72.32%. In [9], the authors combine CNN-based feature extractors with various classifiers, such as SVM and DNN s, achieving an UAR score of 84.97% on the entire SVD dataset.
@@
Finally, in [10], the authors test a combination of SVM and CNN-based feature extractors on spectrograms of sound and electroglottogram (EGG) signals. They integrate these with traditional acoustic features like MFCC, linear predictive coefficients (LPC), and fundamental frequency (f0), reaching an accuracy of 90.10% and UAR we computed as 88.75%. Similarly, they reach an accuracy of 87.41% by extracting the mentioned features and using SVM for classification. However, they reach such high performance without specifying how they handled multiple recordings from the same patient which occur in the dataset, potentially introducing data leakage and therefore reporting an overly optimistic performance.
@@
In [11], the authors take age and sex information from the database and extract various acoustic features from the time domain, such as f0, jitter, shimmer, and HNR. They train ML models, including boosted trees, SVM, decision trees (DT), naive Bayes (NB), and k-nearest neighbors (KNN), achieving the highest accuracy of 84.5% with a boosted tree model on an imbalanced dataset and UAR we computed as 84.55%. In another study [12], they expand their work by incorporating MFCC and their derivatives, and test additional classifiers such as the logistic model tree and instance-based learning algorithms, reaching 85.77% accuracy with a SVM model and UAR we computed as 85.77%. However, they utilized a balanced subset of SVD.
@@
In [13], the authors extract features based on MFCC and self-supervised algorithms from samples of healthy individuals and patients with hyperfunctional dysphonia and vocal fold paresis. Using SVM, they achieve 75.65% and 74.50% accuracy for male and female patients, respectively. Compared to that, [14] focuses on developing classifiers to distinguish between healthy subjects and patients suffering from spasmodic dysphonia and laryngeal nerve paralysis. Their exploration of multi-modal classification methods results in an accuracy of 68.11%.
@@
Further expanding the scope, [15] investigates multi-modal classification for patients with various conditions, including dysphonia, laryngitis, Reinke’s edema, vox senilis, and central laryngeal motion disorder, using energy, zero-crossing rates and entropy as features. By employing SVM and NN, they achieve an average accuracy of 88.46%.
@@
Similarly, [16] explores multimodal classification for conditions such as nodules, polyps, edema, and paralysis, as well as binary classification between healthy and unhealthy individuals. They utilize MFCC as features and a synthetic minority over-sampling technique (SMOTE)-based method for balancing datasets and reach a maximum F1-score of 90% with a CNN-based model for binary classification and we computed UAR as 90%. However, they do not describe how they handle duplicities in SVD dataset and very likely introduce the data leakage due their methodology.
@@
In another study [17], the combination of MFCC and log-mel-frequency spectral coefficients with deep models, using recordings from the SVD and their own database is proposed. This approach results in an accuracy of up to 81.6%. Meanwhile, [18] focuses on patients with dysphonia, chronic laryngitis, and vocal cord paralysis, employing features extracted by a VGGish model in combination with a long short-term memory (LSTM) network. They achieve an F1-score of up to 80% in distinguishing between healthy individuals and those with paralysis.
@@
In [19], the author tests quantitative voice parameters combined with a multi-layered perceptron (MLP) model to classify healthy individuals and those with hyperfunctional dysphonia, laryngitis, and recurrent laryngeal nerve paralysis, achieving 87.5% accuracy.
@@
Additionally, [20] selects 280 samples for pathology detection, extracting MFCC features and using them in a DT model, which results in an accuracy of 67.9%.
@@
Expanding the feature set, [21] uses MFCC, first and second derivatives of MFCC, linear prediction cepstral coefficients, and constant-Q cepstral coefficients with a Bi-LSTM to classify eight selected pathologies, namely dysody, dysphonia, functional dysphonia, hyperfunctional dysphonia, hypofunctional dysphonia, spasmodic dysphonia, vocal cyst polyp, and healthy individuals, achieving an accuracy of 92.7%, which is the only metrics they provide. Finally, [22] examines the impact of MFCC extraction window length on detecting pathologies in patients with dysphonia and reflux laryngitis. Using SVM, they achieve up to 75.1% accuracy and UAR we computed as 75%.
3 Data
@@
Therefore, we use the SVD [28] for our work as it is publicly available, contains a wide range of pathologies, and is the largest of the databases. The SVD was developed by the Phonetics group at the Department of Language Science and Technology, Saarland University, and is available at https://stimmdb.coli.uni-saarland.de/help_en.php4.
@@
For our research, we use the recordings containing the sound /a:/ in neutral pitch for the pathology detection model due to its frequent use in previous studies and clinical protocols [29, 30]. Figure 2 and Figure 3 present examples of healthy and pathological spectrograms, illustrating the /a:/ sound at a neutral pitch from a healthy female subject (patient ID 1) and the same subject with functional dysphonia.
For feature extraction, we consider some limitations to ensure robust and unbiased results in voice pathology detection. The database contains samples from underage patients. Due to developmental changes at a young age, we exclude any recordings of patients under 18 years of age. Research [31, 32] shows that during development, the characteristics of young voices, such as the fundamental frequency, are distinct from those of fully developed voices. These developmental changes could cause problems for the classifier and reduce its ability to accurately detect pathology. By excluding these age groups, we aim to maintain a more consistent and reliable dataset for our analysis.
Another significant issue is the presence of multiple recordings for some patients. For example, the patient with the talker ID 2027 has 24 recordings of the sound / a: / in a neutral pitch. This repetition poses a risk of data leakage, which can affect the results if not addressed properly.
To prevent this risk in a deterministic way, we select the oldest sample of each type by date and recording ID, resulting in a maximum of two recordings per patient, one healthy and one pathological. We believe that this approach minimizes the likelihood that the classifier will learn patient identities, as the patient’s state remains independent of their identity.
In addition, we exclude recordings labeled 1573-a-n.wav and 87-a-n.wav. The former contains two distinct sound recordings, while the latter is corrupted by an artifact, likely caused by hardware or software errors. More over, we exclude several recordings based on the information provided in the comments from the database, stating they contain some artifacts or other problems. Finally, the database contains multiple recordings classified as pathological, while being only singers. We exclude these samples from the study as well.
@@
Table 1: Age distribution among the sex and voice state in the used data
Female Male
@@
4 Feature Extraction
Acoustic features we use can be categorized into time domain, spectral, and cepstral features. We rely on various Python libraries for the feature extraction, namely parselmouth [34] for features related to f0 and formants, spkit [35] for Shannon entropy, torchaudio [36] for linear-frequency cepstral coefficients (LFCC), librosa for the remaining acoustic features, and SciPy [37] and NumpPy [38] to calculate statistical values from the extracted features or from the signal. Table 2 describes libraries used to extract each feature. It also contains the list of all features used in this article, along with references to other related papers utilizing these features. The description of all features are in B.

### v3

@@ Abstract
Purpose: We introduce a novel methodology for voice pathology detection using the publicly available Saarbrücken Voice Database (SVD) and a robust feature set combining commonly used acoustic handcrafted features with two novel ones: pitch difference (relative variation in fundamental frequency) and NaN feature (failed fundamental frequency estimation).
Methods: We evaluate six machine learning (ML) algorithms — support vector machine, k-nearest neighbors, naive Bayes, decision tree, random forest, and AdaBoost — using grid search for feasible hyperparameters and 20480 different feature subsets. Top 1000 classification models – feature subset combinations for each ML algorithm are validated with repeated stratified cross-validation. To address class imbalance, we apply K-Means SMOTE to augment the training data.
Results: Our approach achieves 85.61%, 84.69% and 85.22% unweighted average recall (UAR) for females, males and combined results respectively. We intentionally omit accuracy as it is a highly biased metric for imbalanced data.
Conclusion: Our study demonstrates that by following the proposed methodology and feature engineering, there is a potential in detection of various voice pathologies using ML models applied to the simplest vocal task, a sustained utterance of the vowel /a:/. To enable easier use of our methodology and to support our claims, we provide a publicly available GitHub repository with DOI 10.5281/zenodo.13771573. Finally, we provide a REFORMS checklist to enhance readability, reproducibility and justification of our approach.
keywords:
Voice pathology detection , voice disorder detection , Saarbrücken Voice Database , SVD , machine learning , REFORMS
journal: Journal of Voice
[…]
[1]organization=Department of Mathematics, Informatics, and Cybernetics, University of Chemistry and Technology, Prague, addressline=Technická 5, city=Prague, postcode=166 28, country=Czech Republic
[…]
[2]organization=Department of Radiological Imaging and Informatics, Tohoku University Graduate School of Medicine, addressline=2-1-1 Katahira, Aoba-ku, city=Sendai, postcode=980-8577, country=Japan
[…]
[3]organization=Department of Mathematics and Physics, University of Campania ”Luigi Vanvitelli”, addressline=Viale Abramo Lincoln 5, city=Caserta, postcode=81100, country=Italy
[…]
[4]organization=Department of Otorhinolaryngology, Faculty Hospital Královské Vinohrady, addressline=Šrobárova 1150/50, city=Prague, postcode=100 34, country=Czech Republic
{graphicalabstract}[Uncaptioned image]
@@
Introducing reproducible machine-learning-based system for voice pathology detection.
Novel pitch difference and NaN features improving classification performance.
Addressing class imbalance with k-means SMOTE, boosting minority class predictions.
Avoiding data leakage by managing multiple recordings per patient effectively.
Performing extensive grid search for optimal features and hyperparameters.
@@ 1 Introduction
Voice and speech are fundamental aspects of human communication, playing a crucial role in social interaction, emotional expression, and professional performance. Disorders affecting these functions can have a profound impact on an individual’s quality of life, leading to reduced intelligibility, impaired social interactions, and psychological distress. Voice and speech pathologies encompass a diverse range of conditions, each characterized by distinct alterations in the acoustic and articulatory properties of speech signals. A precise understanding of these pathologies is essential for accurate diagnosis, effective therapeutic intervention, and improved clinical outcomes. [1]
Traditionally, voice disorders have often been described primarily in terms of voice quality disturbances, as seen in conditions such as dysphonia, which is commonly associated with hoarseness, breathiness, or strain. However, this perspective is too narrow to encompass the full spectrum of speech and voice disorders. For instance, dysarthria, a motor speech disorder resulting from neurological impairment, affects articulation, prosody, resonance, and speech duration, rather than merely altering voice quality. Similarly, conditions such as apraxia of speech and neurological voice disorders, including spasmodic dysphonia and hypokinetic dysarthria due to the Parkinson’s disease, lead to a combination of deficits in phonation, speech timing, and intonation. [2, 3] Moreover, emerging research has emphasized the importance of objective acoustic and articulatory analyses in the evaluation of voice and speech disorders.
Advancements in computerized speech analysis, laryngeal imaging, and neurophysiological assessments have provided deeper insights into the pathophysiological mechanisms underlying these conditions. These tools enable clinicians and researchers to characterize speech disorders more precisely, facilitating early diagnosis and the development of targeted interventions.
This involves the processing and analysis of various acoustic features of the voice signal, which can reveal changes or alterations in voice quality caused by specific diseases [4]. In this context, artificial intelligence can be a valid and powerful tool as it may support the decision making process in clinical settings. However, utilization of these solutions imposes many additional tasks, ranging from data collection and preprocessing to ground truth labeling and correct algorithm selection.
@@
In this paper on voice pathology detection, we introduce a methodology for Saarbrücken Voice Database (SVD) that effectively addresses the challenge of multiple recordings from the same patients. Without proper handling, such data redundancy can lead to data leakage, an undesirable effect in which the classification model obtains information about the validation set during the training phase, which might lead to overly optimistic classification results [7]. We consider this an important contribution, as this issue, in the context of voice pathology detection, is rarely discussed in the literature.
Additionally, we present two novel features — pitch difference and NaN feature. Several used features are based on the fundamental frequency (f0). The NaN feature reflects the fact, that we were not able to extract f0 values in the analyzed speech signal. The pitch difference feature quantifies the variability of f0 by measuring fluctuations within a single sustained vowel production. We assume that this variability is often altered in pathological voices due to impairments in vocal fold function, making it a possibly useful feature for voice pathology detection. We experimentally verify proposed features’ usefulness in voice pathology detection.
Based on our experimental findings, we emphasize the necessity of training models separately for male and female patients. Given that we do not investigate the causality or direct correlation of individual features with specific pathologies — and instead work with a large feature set — we leverage machine learning (ML) algorithms. Our feature selection process reveals that the optimal feature sets may vary depending on both the sex and ML algorithm. To identify robust feature sets, we train multiple classification models across different combinations of feature subsets (see Table 2).
By computing the mean Matthews correlation coefficient (MCC), we identify the most promising feature – classification model combinations. We then perform repeated stratified 10-fold cross-validation for the top-performing models to estimate the average performance metrics along with their standard deviations, providing insight into the confidence of our results.
To our best knowledge, our work is the first in the field of voice pathology detection to provide fully reproducible results. The code used for computations and feature extraction is available in our repository, ensuring complete transparency and allowing anyone to verify, build upon, or extend our work. By prioritizing reproducibility, we aim to set a standard for rigorous and open research.
@@ 2 Related Works
In this chapter, we outline the original contribution related to voice pathology detection present in the literature, limiting our research to works connected to SVD. This choice is due to its comprehensiveness: it is the only voice recording database representing the common voice pathologies. See Section 3.1 for detailed information.
@@
Harar et al. [8] test various acoustic features, such as pitch, jitter, shimmer, harmonic-to-noise ratio (HNR), detrended fluctuation analysis parameters, glottis quotients (open, closed), glottal-to-noise excitation ratio, Teager–Kaiser energy operator, modulation energy, and normalized noise energy as well as mel-frequency cepstral coefficients (MFCC). Moreover, they considered the sound samples and their spectrograms as input, all in combination with XGBoost, IsolationForest, and DenseNet models to determine the pathologic samples, reaching an F1 score of 73.3% and unweighted average recall (UAR) we computed as 73.3%. While, in Gupta et al.[9], features derived from self-supervised learning models, Data2Vec and Wav2Vec, alongside MFCC are explored. The reliability of these features to evaluate voice quality is tested using support vector machine (SVM) and deep neural networks (DNN), achieving an accuracy of 77.83% and UAR we computed as 77.86%.
@@
A different approach is taken by Verde et al.[10], where the authors transform sound wave data into spectrograms and treat classification as an image recognition problem using a convolutional neural network (CNN), achieving 73.93% accuracy and UAR we computed as 70.68%.
@@
In Kotarba & Kotarba[11], the authors use MFCC and gammatone frequency cepstral coefficients (GFCC) with a classification model based on a neural network (NN), achieving 81.84% accuracy. Unlike other studies that use the /a:/ sound for feature extraction, they employ whole sentences. Unfortunately, we were not able to compute UAR, because the reported results appear inconsistent, making it impossible to calculate the metric. Another study by Harar et al.[12] tests a DNN with convolutional layers, using the voice signal as a feature set for the convolutional layer, achieving 68.08% accuracy and UAR we computed as 72.32%. Park et al.[13] combine CNN-based feature extractors with various glsml algorithms, such as SVM and DNN s, achieving an UAR score of 84.97% on the entire SVD dataset.
@@
Finally, Omeroglu et al.[14] test a combination of SVM and CNN-based feature extractors on spectrograms of sound and electroglottogram (EGG) signals. They integrate these with traditional acoustic features like MFCC, linear predictive coefficients (LPC), and f0, reaching an accuracy of 90.10% and UAR we computed as 88.75%. Similarly, they reach an accuracy of 87.41% by extracting the mentioned features and using SVM for classification. However, they reach such high performance without specifying how they handled multiple recordings from the same patient which occur in the dataset, potentially introducing data leakage and therefore reporting an overly optimistic performance.
@@
In another study by Verde et al.[15], the authors take age and sex information from the database and extract various acoustic features from the time domain, such as f0, jitter, shimmer, and HNR. They employ ML algorithms, including boosted trees, SVM, decision trees (DT), naive Bayes (NB), and k-nearest neighbors (KNN), achieving the highest accuracy of 84.5% with a boosted tree model on an imbalanced dataset and UAR we computed as 84.55%. In the follow-up study [16], they expand their work by incorporating MFCC and their derivatives, and test additional ML algorithms such as logistic model tree and instance-based learning algorithms, reaching 85.77% accuracy with a SVM model and UAR we computed as 85.77%. However, they utilized a balanced subset of SVD.
@@
Tirronen et al.[17] extract features based on MFCC and self-supervised algorithms from samples of healthy individuals and patients with hyperfunctional dysphonia and vocal fold paresis. Using SVM, they achieve 75.65% and 74.50% accuracy for male and female patients, respectively. Compared to that, Yagnavajjula et al.[18] focus on developing classification models to distinguish between healthy subjects and patients suffering from spasmodic dysphonia and laryngeal nerve paralysis. Their exploration of multi-modal classification methods results in an accuracy of 68.11%.
@@
Further expanding the scope, Junior et al.[19] investigate multi-modal classification for patients with various conditions, including dysphonia, laryngitis, Reinke’s edema, vox senilis, and central laryngeal motion disorder, using energy, zero-crossing rate, and entropy as features. By employing SVM and NN, they achieve an average accuracy of 88.46%.
@@
Similarly, Fan et al.[20] explore multimodal classification for conditions such as nodules, polyps, edema, and paralysis, as well as binary classification between healthy and unhealthy individuals. They utilize MFCC as features and a synthetic minority over-sampling technique (SMOTE)-based method for balancing datasets and reach a maximum F1-score of 90% with a CNN-based model for binary classification and we computed UAR as 90%. However, they do not describe how they handle duplicities in SVD dataset and very likely introduce the data leakage due their methodology.
@@
Ding et al.[21] propose their own approach by combining MFCC and log-mel-frequency spectral coefficients with deep models, using recordings from the SVD and their own database. This approach results in an accuracy of up to 81.6%. Meanwhile, Guedes et al.[22] focus on patients with dysphonia, chronic laryngitis, and vocal cord paralysis, employing features extracted by a VGGish model in combination with a long short-term memory (LSTM) network. They achieve an F1-score of up to 80% in distinguishing between healthy individuals and those with paralysis.
@@
Hemmerling[23] tests quantitative voice parameters combined with a multi-layered perceptron (MLP) model to classify healthy individuals and those with hyperfunctional dysphonia, laryngitis, and recurrent laryngeal nerve paralysis, achieving 87.5% accuracy.
@@
Additionally, AL-Dhief et al.[24] select 280 samples for pathology detection, extracting MFCC features and using them in a DT model, which results in an accuracy of 67.9%.
@@
Expanding the feature set, AnilKumar & Reddy[25] use MFCC, first and second derivatives of MFCC, linear prediction cepstral coefficients, and constant - Q cepstral coefficients with a Bi-LSTM to classify eight selected pathologies, namely dysody, dysphonia, functional dysphonia, hyperfunctional dysphonia, hypofunctional dysphonia, spasmodic dysphonia, vocal cyst polyp, and healthy individuals, achieving an accuracy of 92.7%, which is the only metric they provide. Finally, Tirronen et al.[26] examine the impact of MFCC extraction window length on detecting pathologies in patients with dysphonia and reflux laryngitis. Using SVM, they achieve up to 75.1% accuracy and UAR we computed as 75%.
Finally, multiple works study the use of end-to-end models, which are not solely dependent on hand-crafted features. [27, 28] Liu et al. [29] propose an end-to-end deep learning model for classification of laryngitis and hyperfunctional dysphonia using stacked vowels, while reaching UAR of 72%. Reddy et al. [30] apply a wavelet scattering network to extract features from recordings in combination with MLP for classification. With their end-to-end approach, they reach 81.32% UAR using recordings of sentences as their input.
3 Materials & Methods
To make our methodology more comprehensive, we provide a flowchart to show the process of feature extraction, training and results validation in Figure 1, along with the reference to the relevant tables and figures in the manuscript.
Figure 1: Flow diagram of the proposed methodology
3.1 Data
@@
Therefore, we use the SVD [36] for our work as it is publicly available, contains a wide range of pathologies, and is the largest of the databases. The SVD was developed by the Phonetics group at the Department of Language Science and Technology, Saarland University, and is available at https://stimmdb.coli.uni-saarland.de.
@@
For our research, we use the recordings containing the sound /a:/ in neutral pitch for pathology detection due to its frequent use in previous studies and clinical protocols [37, 38]. Figure 2 and Figure 3 present examples of healthy and pathological spectrograms, illustrating the /a:/ sound in neutral pitch from a healthy female subject (patient ID 1) and the same subject with functional dysphonia. We fully recognize that many voice as well as speech pathologies go beyond phonatory function and involve articulatory, prosodic, and temporal characteristics, which are more effectively captured in continuous speech.
For feature extraction, we consider some limitations to ensure robust and unbiased results in voice pathology detection. The database contains samples from underage patients. Due to developmental changes at a young age, we exclude any recordings of patients under 18 years of age. Research [39, 40] shows that during development, the characteristics of young voices, such as the fundamental frequency, are distinct from those of fully developed voices. These developmental changes could cause problems for the classification model and reduce its ability to accurately detect pathology. By excluding these age groups, we aim to maintain a more consistent and reliable dataset for our analysis.
Another significant issue is the presence of multiple recordings for some patients. For example, the patient with the talker ID 2027 has 24 recordings of the sound /a:/ in neutral pitch. This repetition poses a risk of data leakage, which can affect the results if not addressed properly.
To prevent this risk in a deterministic way, we select the oldest sample of each type by date and recording ID, resulting in a maximum of two recordings per patient, one healthy and one pathological. We believe that this approach minimizes the likelihood that the classification model learns patient identities, as the patient’s state remains independent of their identity.
In addition, we exclude recordings labeled 1573-a-n.wav and 87-a-n.wav. The former contains two distinct sound recordings, while the latter is corrupted by an artifact, likely caused by hardware or software errors. Moreover, we exclude several recordings based on the information provided in the comments from the database, stating they contain some artifacts or other problems. Finally, for several recordings, the ”Pathologies” column contains information that the speaker performs the tasks during the recording using their singing voice or that they are a singer, which may lead to incorrect labeling of the subject as pathological. We exclude these samples from the study as well.
@@
Table 1: Age distribution among the sex and voice condition in the used data
Female Male Unit
@@
3.2 Feature Extraction
Acoustic features we use can be categorized into time domain, spectral, and cepstral features. We rely on various Python libraries for the feature extraction, namely parselmouth [42] for features related to f0 and formants, spkit [43] for Shannon entropy, torchaudio [44] for linear-frequency cepstral coefficients (LFCC), librosa for the remaining acoustic features, and SciPy [45] and NumpPy [46] to calculate statistical values from the extracted features or from the signal. Table 2 describes libraries used to extract each feature. It also contains the list of all features used in this article, along with references to other related papers utilizing these features.